# OODTE: A Differential Testing Engine for the ONNX Optimizer

### v3

@@ 6.1. Optimizer Crashes
Out of the $15$ detected bugs, $12$ bugs were related to model crashes, $2$ bugs resulted in output differences between the original and the optimized model, and $1$ bug introduced redundant initializers in the model that produced runtime warnings. In particular, we observed crashes were caused by $8$ out of the $47$ passes (1717%), and occurred for specific models, with object detection models being more prone to errors from these $8$ passes. In particular, R-CNN models, such as Faster R-CNN R-50-FPN (faster-rcnn, ), RetinaNet (retinanet, ), VGG-16, VGG-19 (vgg, ), as well as YOLOv3 (Redmon2018YOLOv3AI, ), presented crashes by using the optimizer, in a range of opset instances and in passes related to the modification of gather nodes, and the elimination of idempotent operations. In total, we observed $31$ model instances where either the optimizer crashed upon processing them, or the optimized model was problematic, leading to a runtime crash. Some secondary passes are known to be unstable (e.g., split_init and split_predict describe in their documentation that they not strictly comply with the ONNX format), and therefore we do not focus on them as sources of faults in the optimization process. Overall, $12$ out of the $31$ model instances (9.29.2%) that were associated with a crash occurred while using passes related to fuse and eliminate operations, which are part of the default bundle of optimizations in the ONNX Optimizer and are therefore of particular significance.
@@
Table 2. Overview of the issues identified using OODTE, detailing the responsible optimization pass, the detected error, its type (Crash, Accuracy (Acc.), or Warning (Warn.)), and the models in which the error was observed, along with their opsets (op.).
@@
1 lift_lexical_references
Removes used
value references.
Crash
EfficientNet-Lite4 (op. 11), SSD (op. 12)
2 split_init
List index
out of range.
Crash YOLOv3 (op. 10, 12)
3
split_init
split_predict
Required Field
is missing.
Crash
DenseNet121 (op. 7-9), YOLOv2 (op. 10),
SSD (op. 12), Tiny-YOLOv3 (op. 11),
YOLOv3 (op. 10, 12)
4 split_predict
Required inputs
are missing.
Crash
SSD (op. 12), DenseNet121 (op. 7),
InceptionV2 (op. 6-9)
5 split_predict Bad spec for node. Crash
YOLOv3 (op. 10), DenseNet121 (op. 7),
InceptionV2 (op. 6-9)
6 split_predict
Mismatch in number
of input elements
Crash
FCN ResNet-50-int8 (op. 12)
7 eliminate_shape_gather
Indices value
not equal to
node dims size
Crash
Faster R-CNN: R-50-FPN (op. 10)
8 All Passes
Implicit ONNX
format version
update.
Acc.
Models of cases 11-$11$
in Table 3
9 rewrite_input_dtype
Output name
used multiple times.
Crash GPT-2 (op. 10)
10
eliminate_consecutive_idempotent_ops
Incompatible shape
error.
Crash
Mask R-CNN/Faster R-CNN: R-50-FPN-int8 (12),
R-50-FPN-qdq (op. 12), R-50-FPN-fp32 (op. 12)
11
eliminate_duplicate_initializer
Erase output
assertion failed.
Crash
Faster R-CNN: R-50-FPN-qdq (12)
12 fuse_bn_into_conv
Duplicate initializer
introduced.
Crash
RetinaNet (op. 9)
13
eliminate_shape_gather
eliminate_nop_reshape
Added initializers
not used by the model.
Warn.
Faster R-CNN: R-50-FPN-int8 (op. 12)
14 All Passes Invalid input scale. Crash
VGG-16-BN (op. 7), VGG-19-BN (op. 7),
ResNet101_DUC_HDC (op. 7)
15 fuse_bn_into_conv Bad graph structure. Acc.
Efficient-Lite4 (op. 11),
SSD (op. 12)
@@ 6.2. Output Label Differences in Optimized Classification & Object Detection Models
Out of the $130$ optimized models tested, we observed unexpected output differences between the original and optimized models. These differences manifested as (1) label differences for classification models (e.g., ResNet), (2) label and/or bounding box differences for object detection models (e.g., YOLOv3), and (3) textual output differences in token sequences or predicted spans for text generation tasks. Since the ONNX Optimizer transformations are graph-based and do not alter numeric precision, output differences are not expected. Among the $93$ classification models tested, $28$ (30.130.1%) exhibited discrepancies in output labels. Similarly, $5$ out of $30$ (16.616.6%) object detection models presented output differences. The text generation models, on the other hand, were robust with only one instance (GPT-2) exhibiting negligible differences ($0.001\%$) while all others showing identical results across the original and the optimized model versions. The differences impacted the top-5 and top-10 predictions, while top-1 predictions remained unchanged. Although top-1 accuracy is often emphasized, variations in lower-ranked predictions can also reveal important issues. Our results show that focusing exclusively on top-1 accuracy may give developers a false sense that an optimization is successful, while overlooking substantial changes in top-K predictions (up to 2% in top-5 and 13% in top-10). Using OODTE, we detected these inconsistencies and reported them to the ONNX Optimizer developers.
@@
1 AlexNet (alexnet, ) CL 7-9 0% 0% 0.02% FV
2 DenseNet121 (densenet, ) CL 7-9 0% 0% 0.02% FV
3 GoogleNet (googlenet, ) CL 7-9 0% 0% 0.04% FV
4 Inception-1 (inception, ) CL 7-9 0% 0% 0.25% FV
5 Inception-2 (inception, ) CL 7-9 0% 0.02% 0.05% FV
6 MobileNet v2 (mobilenet, ; mobilenetv2, ) CL 7-9 0% 0% 0.02% FV
7 ResNet18 (resnet, ) CL 7 0% 0.02% 0.02% FV
8 ResNet34-v2 (resnetv2, ) CL 7 0% 0.02% 0.02% FV
9 ResNet50-caffe2 (resnet, ) CL 7-9 0% 2% 13% FV
10 ShuffleNet-v1 (shufflenet, ) CL 7-9 0% 0.7% 1% FV
11 SqueezeNet 1.0 (squeezenet, ) CL 7-9 0% 0% 0.02% FV
12 EfficientNet-Lite4 (efficientnet, ) CL 11 0% 0.04% 1% MG
13 SSD (ssd, ) OD 12 0% 0% 0.04% MG
@@ 9 ResNet50-caffe2 (resnet, ) CL 7-9 0% 2% 13% FV
fuses Batch Normalization and Convolution operators within the model. In particular, the pass attempts to fuse BatchNorma-
lization
@@
SSD (ssd, ) 12 99.99% 99.99% 99.99% 99.99% 100%
Tiny
YOLOv3 (Redmon2018YOLOv3AI, )
11 99.18% 99.18% 99.18% 99.99% 89.35%
YOLOv2 (yolov2, ) 9 99.95% 99.95% 99.95% 99.94% 88.49%
YOLOv3 (Redmon2018YOLOv3AI, ) 10 99.45% 99.45% 99.45% 99.99% 84.81%
YOLOv3 (Redmon2018YOLOv3AI, ) 12 99.45% 99.45% 99.45% 99.99% 84.81%
@@
Bounding Box Differences in Object Detection Models We observed some differences in the bounding box outputs between the original and optimized versions of object detection models. The results are presented in detail in Table 4 with the following main findings: 1. Both YOLOv3 models (opset $10$ and $12$) achieve high average IoU, F1 score, precision, and recall, all above 99.4599.45%. However, the optimized model occasionally misplaces bounding boxes—resulting in IoU drops of up to 15.1915.19%, these differences do not lead to misclassifications, as none were observed in the top ranked label predictions. A similar trend is also seen in YOLOv2 model where optimization causes a drop in IoU value but not in detection performance (F1, precision and recall).
@@ 7. Tool Usability
OODTE  can test whether the ONNX Optimizer generates optimized models without compromising output label correctness. Currently, OODTE supports differential testing for classification, object detection, and text generation models. However, OODTE can also be easily extended to support new model types, as long as (1) the models can be compiled into the ONNX format for optimization and (2) a comparator is provided to compare the results between the base and optimized models for the new model type. In addition, OODTE can fetch models and datasets automatically, making it adaptable for use in a Continuous Integration pipeline. This allows OODTE to test the ONNX Optimizer whenever new functionality (e.g., new passes) is added. By default, OODTE loads models from the official ONNX Model Hub (onnxmodelhub, ), which was used in our experiments. However, by configuring OODTE to use a custom ONNX Hub that follows the ONNX API specification for model repositories (onnxhubapi, ), it can fetch and use models without further modification. The same flexibility applies to datasets. OODTE can fetch datasets using packages like the datasets (datasets, ) package or from local storage.
Finally, while OODTE applies a differential testing methodology focused on the ONNX Optimizer, this approach is generalizable and can be easily adapted to other AI optimizers or compilers that meet two primary requirements: (1) they provide a runtime for model execution, and (2) they expose an API that includes all the passes applied during model optimization, both as a bundle and as independent passes.
@@ 8. Threats To Validity
Methodology Generalizability: We focus ONNX Optimizer because it is a core component of the widely used ONNX ecosystem (19.5K stars) (onnx, ) and supports tasks beyond model optimization, such as graph simplification (ONNX Simplifier, 4.1K stars (onnxsimplifier, )). Its popularity and broad applicability make it a strong candidate for study and a representative of the state-of-the-art technologies. While testing all DL compilers was infeasible, we demonstrated the value of our methodology and noted it can extend to other optimizers that offer per-pass APIs and runtime support for differential testing. For example, OODTE could support OpenVINO (openvino, ) and Apache TVM (tvm, ) by adapting their graph optimization passes, considering both compilers also provide runtime APIs, similarly to ONNX.
@@
The order of optimization passes can significantly affect model performance and behavior (OptimizationOrder, ), a well-known challenge in compiler research. While we do not study pass ordering in depth, OODTE adopts two evaluation modes: (1) the default ONNX Optimizer configuration, including common passes such as fuse and eliminate, and (2) a configuration applying all 47 available passes atomically. Although this does not capture issues from specific pass combinations or orderings, it effectively uncovers correctness problems in practice. Nevertheless, the omission of a systematic study on optimization ordering may limit the completeness of our evaluation.
@@ 9. Discussion
We summarize below our findings for the different model types in our experiments.
(1) Optimizations can affect label accuracy in classification models. We identified two primary causes for label accuracy differences after optimization: (a) Legacy opset upgrade: Some models used older ONNX opsets, and the optimizer implicitly and improperly upgraded their format versions. Classification models built on these legacy opsets were particularly affected. (b) Problematic graph node generation: The fuse_bn_into_conv pass, which fuses batch normalization into convolutional layers, can unintentionally alter model behavior and particularly affects classification models that rely on convolutions.
Although text generation and comprehension models are also large, manual inspection showed that the optimizer applies only a few passes to them. Most optimizations are skipped because the targeted layer patterns (e.g., convolutions) are uncommon in transformer-based architectures used for language tasks.
(2) Object detection models crash more frequently when optimized.
Most crashes during optimization occurred in object detection and semantic segmentation models, rather than in classification models. This is likely due to their greater architectural complexity—being larger and structurally more complex—which leads the ONNX Optimizer to apply more extensive transformations, increasing the likelihood of encountering corner cases that trigger crashes related to the ONNX Optimizer passes.
(3) Optimization can affect bounding boxes of object detection models. Regarding accuracy, object detection models remained largely stable. We observed only a negligible decrease in the accuracy of SSD on top-10 accuracy predictions (0.040.04%), while the other models showed no difference between the original and optimized versions. However, the optimized models occasionally misaligned bounding boxes, leading to drops in IoU value of up to 15.1915.19%.
(4) Outputs from text comprehension and generation models were not affected by optimization. Among the 8 text comprehension and generation models tested, we observed only one crash (in GPT-2) and a small number of samples (8) with BLEU-score differences between the original and optimized models for the top-3 token selections.
Overall, we find ONNX optimizations can result in output differences between the optimized and unoptimized model, with some model types being more susceptible to optimization errors than others. OODTE was able to successfully isolate the optimization passes responsible for the output differences for different model types.

### v5

@@ 6.1. Optimizer Crashes
Out of the $15$ detected bugs, $12$ bugs were related to model crashes, $2$ bugs resulted in output differences between the original and the optimized model, and $1$ bug introduced redundant initializers in the model that produced runtime warnings. In particular, we observed crashes were caused by $8$ out of the $47$ passes (1717%), and occurred for specific models, with object detection models being more prone to errors from these $8$ passes. In particular, R-CNN models, such as Faster R-CNN R-50-FPN (Ren et al., 2015), RetinaNet (Lin et al., 2020), VGG-16, VGG-19 (Simonyan and Zisserman, 2015), as well as YOLOv3 (Redmon and Farhadi, 2018), presented crashes by using the optimizer, in a range of opset instances and in passes related to the modification of gather nodes, and the elimination of idempotent operations. In total, we observed $31$ model instances where either the optimizer crashed upon processing them, or the optimized model was problematic, leading to a runtime crash. Some secondary passes are known to be unstable (e.g., split_init and split_predict describe in their documentation that they not strictly comply with the ONNX format), and therefore we do not focus on them as sources of faults in the optimization process. Overall, $12$ out of the $31$ model instances (9.29.2%) that were associated with a crash occurred while using passes related to fuse and eliminate operations, which are part of the default bundle of optimizations in the ONNX Optimizer and are therefore of particular significance.
@@
Table 2. Overview of the issues identified using DiTOX, detailing the responsible optimization pass, the detected error, its type (Crash, Accuracy (Acc.), or Warning (Warn.)), and the models in which the error was observed, along with their opsets (op.). The (Run) label in crashes indicates that the model was successfully optimized but failed during execution.
@@
1 lift_lexical_references Removes used value references. Crash EfficientNet-Lite4 (op. 11), SSD (op. 12)
2 split_init Out of range List index Crash YOLOv3 (op. 10, 12)
3 split_init split_predict Missing required field Crash DenseNet121 (op. 7-9), YOLOv2 (op. 10), SSD (op. 12), Tiny-YOLOv3 (op. 11), YOLOv3 (op. 10, 12)
4 split_predict Missing required inputs Crash SSD (op. 12), DenseNet121 (op. 7), InceptionV2 (op. 6-9)
5 split_predict Bad spec for node. Crash YOLOv3 (op. 10), DenseNet121 (op. 7), InceptionV2 (op. 6-9)
6 split_predict Mismatch in #input elements Crash FCN ResNet-50-int8 (op. 12)
7 eliminate_shape_gather Indices value $\neq$ node dims size. Crash Faster R-CNN: R-50-FPN (op. 10)
8 All Passes Implicit ONNX format version update. Acc. Models of cases 11-$11$ in Table 3
9 rewrite_input_dtype Multiple usage of output name Crash GPT-2 (op. 10)
10 eliminate_consecutive _idempotent_ops Incompatible shape error. Crash (Run) Mask R-CNN/Faster R-CNN: R-50-FPN-int8 (12), R-50-FPN-qdq (op. 12), R-50-FPN-fp32 (op. 12)
11 eliminate_duplicate_initializer Erase output assertion failed. Crash Faster R-CNN: R-50-FPN-qdq (12)
12 fuse_bn_into_conv Duplicate initializer introduced. Crash (Run) RetinaNet (op. 9)
13 eliminate_shape_gather eliminate_nop_reshape Unused initializers Warn. Faster R-CNN: R-50-FPN-int8 (op. 12)
14 All Passes Invalid input scale. Crash VGG-16-BN (op. 7), VGG-19-BN (op. 7), ResNet101_DUC_HDC (op. 7)
15 fuse_bn_into_conv Bad graph structure. Acc. Efficient-Lite4 (op. 11), SSD (op. 12)
@@ 6.2. Output Label Differences in Optimized Classification & Object Detection Models
Out of the $130$ optimized models tested, we observed unexpected output differences between the original and optimized models. These differences manifested as (1) label differences for classification models (e.g., ResNet), (2) label and/or bounding box differences for object detection models (e.g., YOLOv3), and (3) textual output differences in token sequences or predicted spans for text generation tasks. Since the ONNX Optimizer transformations are graph-based and do not alter numeric precision, output differences are not expected. Among the $93$ classification models tested, $28$ (30.130.1%) exhibited discrepancies in output labels. Similarly, $5$ out of $30$ (16.616.6%) object detection models presented output differences. The text generation models, on the other hand, were robust with only one instance (GPT-2) exhibiting negligible differences ($0.001\%$) while all others showing identical results across the original and the optimized model versions. The differences impacted the top-5 and top-10 predictions, while top-1 predictions remained unchanged. Although top-1 accuracy is often emphasized, variations in lower-ranked predictions can also reveal important issues. Our results show that focusing exclusively on top-1 accuracy may give developers a false sense that an optimization is successful, while overlooking substantial changes in top-K predictions (up to 2% in top-5 and 13% in top-10). Using DiTOX, we detected these inconsistencies and reported them to the ONNX Optimizer developers.
@@
1 AlexNet (Krizhevsky et al., 2017) CL 7-9 0% 0% 0.02% FV
2 DenseNet121 (Huang et al., 2018) CL 7-9 0% 0% 0.02% FV
3 GoogleNet (Szegedy et al., 2014a) CL 7-9 0% 0% 0.04% FV
4 InceptionV1 (Szegedy et al., 2014b) CL 7-9 0% 0% 0.25% FV
5 InceptionV2 (Szegedy et al., 2014b) CL 7-9 0% 0.02% 0.05% FV
6 MobileNetV2 (Howard et al., 2017; Sandler et al., 2018) CL 7-9 0% 0% 0.02% FV
7 ResNet18 (He et al., 2015) CL 7 0% 0.02% 0.02% FV
8 ResNet34V2 (He et al., 2016) CL 7 0% 0.02% 0.02% FV
9 ResNet50-caffe2 (He et al., 2015) CL 7-9 0% 2% 13% FV
10 ShuffleNetV1 (Zhang et al., 2017) CL 7-9 0% 0.7% 1% FV
11 SqueezeNet 1.0 (Iandola et al., 2017) CL 7-9 0% 0% 0.02% FV
12 EfficientNet-Lite4 (10) CL 11 0% 0.04% 1% MG
13 SSD (Liu et al., 2016) OD 12 0% 0% 0.04% MG
@@
fuses Batch Normalization and Convolution operators within the model. In particular, the pass attempts to fuse BatchNor-
malization
@@
SSD (Liu et al., 2016) 12 99.99% 99.99% 99.99% 99.99% 100%
Tiny YOLOv3 (Redmon and Farhadi, 2018) 11 99.18% 99.18% 99.18% 99.99% 89.35%
YOLOv2 (Redmon and Farhadi, 2016) 9 99.95% 99.95% 99.95% 99.94% 88.49%
YOLOv3 (Redmon and Farhadi, 2018) 10 99.45% 99.45% 99.45% 99.99% 84.81%
YOLOv3 (Redmon and Farhadi, 2018) 12 99.45% 99.45% 99.45% 99.99% 84.81%
@@
Bounding Box Differences in Object Detection Models We observed some differences in the bounding box outputs between the original and optimized versions of object detection models. The results are presented in detail in Table 4 with the following main findings: 1. Both YOLOv3 models (opset $10$ and $12$) achieve high average IoU, F1 score, precision, and recall, all above 99.4599.45%. However, the optimized model occasionally misplaces bounding boxes—resulting in IoU drops of up to 15.1915.19%, these differences do not lead to misclassifications, as none were observed in the top ranked label predictions. A similar trend is also seen in YOLOv2 model where optimization causes a drop in IoU value but not in detection performance (F1, precision and recall). Although this behavior may appear minor, it can lead to undesirable consequences, particularly in safety-critical systems where even small deviations matter. While our experiments did not observe bounding box misclassifications, such outcomes cannot be ruled out for other models or different inputs.
@@ 7. Tool Usability
DiTOX  can test whether the ONNX Optimizer generates optimized models without compromising output label correctness. Currently, DiTOX supports differential testing for classification, object detection, and text generation models. However, DiTOX can also be easily extended to support new model types, as long as (1) the models can be compiled into the ONNX format for optimization and (2) a comparator is provided to compare the results between the base and optimized models for the new model type. In addition, DiTOX can fetch models and datasets automatically, making it adaptable for use in a Continuous Integration pipeline. This allows DiTOX to test the ONNX Optimizer whenever new functionality (e.g., new passes) is added. By default, DiTOX loads models from the official ONNX Model Hub (39), which was used in our experiments. However, by configuring DiTOX to use a custom ONNX Hub that follows the ONNX API specification for model repositories (38), it can fetch and use models without further modification. The same flexibility applies to datasets. DiTOX can fetch datasets using packages like the datasets (8) package or from local storage.
8. Tool Generalizability
Although DiTOX applies differential testing to the ONNX Optimizer, the approach is generic and can be adapted to other AI optimizers or compilers that provide a runtime for model execution and expose optimization passes both collectively and individually through an API. Many compiler ecosystems already satisfy these requirements. For example, MLIR (Lattner et al., 2020b) offers passes such as graph rewrites, Apache TVM (Chen et al., 2018) provides a Sequential API for composing optimizations like constant folding and subexpression elimination, and OpenVINO (Zunin, 2021) supports a full set of individual passes. While we focus on ONNX due to its wide adoption, the methodology of DiTOXis model-format independent and can be applied to other frameworks using their supported model types and runtimes. Finally, DiTOX includes comparators for classifiers, object detectors, and transformers; for other model families, users can implement custom comparators by following the already provided comparator API.
@@ 9. Threats To Validity
Methodology Generalizability: We focus ONNX Optimizer because it is a core component of the widely used ONNX ecosystem (19.5K stars) (43) and supports tasks beyond model optimization, such as graph simplification (ONNX Simplifier, 4.1K stars (42)). Its popularity and broad applicability make it a strong candidate for study and a representative of the state-of-the-art technologies. While testing all DL compilers was infeasible, we demonstrated the value of our methodology and noted it can extend to other optimizers that offer per-pass APIs and runtime support for differential testing. For example, DiTOX could support OpenVINO (Zunin, 2021) and Apache TVM (Chen et al., 2018) by adapting their graph optimization passes, considering both compilers also provide runtime APIs, similarly to ONNX.
@@
The order of optimization passes can significantly affect model performance and behavior (Kulkarni and Cavazos, 2012), a well-known challenge in compiler research. While we do not study pass ordering in depth, DiTOX adopts two evaluation modes: (1) the default ONNX Optimizer configuration, including common passes such as fuse and eliminate, and (2) a configuration applying all 47 available passes atomically. Although this does not capture issues from specific pass combinations or orderings, it effectively uncovers correctness problems in practice. Nevertheless, the omission of a systematic study on optimization ordering may limit the completeness of our evaluation.
@@ 10. Discussion
We summarize our findings across different model types below.
(1) Optimizations can affect label accuracy in classification models. Label discrepancies in classification models arose primarily from two sources. First, models using legacy ONNX opsets were implicitly and incorrectly upgraded by the optimizer, disproportionately affecting classification models. Second, the fuse_bn_into_conv pass occasionally generated problematic graph nodes, altering model behavior in convolution-heavy architectures. In contrast, text models were largely unaffected, as the optimizer applies few passes to transformer-based architectures where convolutional patterns are rare.
(2) Object detection models are more prone to optimization crashes. Most optimizer crashes occurred in object detection and semantic segmentation models. Their larger size and structural complexity trigger more aggressive transformations, increasing exposure to corner cases in optimization passes.
(3) Optimization can degrade bounding box accuracy. While overall detection accuracy remained stable, optimized models occasionally produced misaligned bounding boxes. This resulted in IoU drops of up to 15.1915.19%, despite negligible changes in top-10 accuracy (e.g., 0.040.04% for SSD).
(4) Text models are largely robust to optimization. Among the eight text comprehension and generation models, we observed only one optimizer crash (GPT-2) and minor BLEU-score differences affecting eight samples in top-3 token predictions.
Overall, ONNX optimizations can introduce correctness issues, with susceptibility varying by model type. DiTOX effectively isolates the optimization passes responsible for these discrepancies across model classes.
11. Future Work
We identify three potential extensions of this work as future research directions:
(1) Execution Time Measurements: This study primarily focused on optimizer correctness, specifically on detecting crashes and accuracy faults. However, in the broader context of model optimization, it is also important to assess how an optimizer impacts overall model performance. In particular, this includes execution time for both CPU and GPU deployments, as well as other factors such as model size. DiTOX can be easily extended to support hardware-accelerated model execution, either by leveraging the ONNX Runtime API (41) or by integrating a third-party AI compiler that generates target-specific device code, such as Apache TVM (Chen et al., 2018). We leave this direction for future work, as it could provide valuable additional insights.
(2) Selection of Different Model Types: This work focused on classification, object detection, and text generation models. However, DiTOX is applicable to a broader range of model types, including speech recognition, super-resolution, and body recognition. We leave the exploration of these additional model categories as a direction for future work.
(3) Examination of the Optimization Pass Order:
DiTOX supports optimization using a user-defined subset and ordering of passes through configuration. Although such experiments could potentially reveal additional faults, we exclude them for tractability, as exhaustively generating, executing, and evaluating all pass permutations is computationally infeasible. We leave this direction for future work, potentially in combination with techniques for reducing the permutation search space.